Suppose the oil price in world market is $60 per gallon,the u.s domestic demand curve of oil is Qd=200-2q and a domestic supply curve is Qs=p
a) if u.s government imposes a tariff of$15 per gallon,what will be the US price amd the level of import?How much revenue will the government earn from tariff?How large is the dead weight loss?
b)if the US government has no tariff but imposes an import quota of 25,what will be the domestic price?what is the cost of the quota for US consumers of the oil?what is the gain for US producers?
1
Expert's answer
2018-02-26T09:11:07-0500
Equilibrium: Qs=Qd => 200-2p=p, =>p=66.67, Q=66.67
a) If u.s government imposes a tariff of$15 per gallon, the world price for US becomes 60+15=$75, 75>66.67, all the demand is satisfied in the domestic market. TR=0
Deadweight loss is ½*(66.67-60)*(200-2*60-60)=16.7


b) Before imposing import quota import was 200-2*60-60=20. Import quota of 25 does not change anything, because 25>20